OpenStudy (anonymous):

simplify the expression: x+8/x^2-5x-6 + x+1/x^2-4x-5

OpenStudy (anonymous):

damn

OpenStudy (whpalmer4):

Is that supposed to be\[\frac{x+8}{x^2-5x-6} + \frac{x+1}{x^2-4x-5}\]by any chance?

OpenStudy (whpalmer4):

you really need to put parentheses around the numerator and denominator if you're going to write it all on one line like that: (x+8)/(x^2-5x-6) + (x+1)/(x^2-4x-5) What you've written is really \[x + \frac{8}{x^2}-5x-6 + x + \frac{1}{x^2}-4x-5\]but I'm confident that isn't what you intend

OpenStudy (whpalmer4):

Anyhow, to simplify this, you'll want to factor both denominators, then get a common denominator so you can combine the fractions

OpenStudy (anonymous):

hey whpalmer4, can you help me?

OpenStudy (whpalmer4):

is that not what I'm doing?

OpenStudy (anonymous):

oh no that's someone eles question. mine is: One railroad track was laid out on the line y = 8x - 2 by the engineers in charge. Which of the following lines could have been the other parallel track? A. y = 2x - 8 B. y = -⅛x + 3 C. y = 4 + 8x D. 2y = 8x + 3

OpenStudy (anonymous):

thanks @whpalmer4

OpenStudy (whpalmer4):

@sagek54321 very rude to inject your question in someone else's question like that. Tag me on your question and I'll have a look when I get a minute

OpenStudy (anonymous):

Im sorry I didnt mean to be rude i just thought that you could help me here.. im new to this site

OpenStudy (anonymous):

@whpalmer4 yes it is suppoe to look like this: \[\frac{ x+8 }{ x^2-5x-6 }+\frac{ x+1 }{ x^2-4x-5 }\]

OpenStudy (whpalmer4):

no problem, I told you what the etiquette is, now it's up to you :-) to tag me, just put @whpalmer4 in the body of a response to your question

OpenStudy (whpalmer4):

okay, can you factor the two denominators?

OpenStudy (anonymous):

yes I can

OpenStudy (anonymous):

So I got this when I factored the denominators.. would the x-5 on the left side and right side both cancel each other? \[\frac{ x+8 }{(x-1)(x-5) }+\frac{ x+1 }{ (x+1)(x-5) }\]

OpenStudy (whpalmer4):

That's not quite a correct factoring, I'm afraid... \[(x-1)(x-5) = x^2 - 5x - 1x + 5 = x^2 -6x + 5\]which is not the denominator of either fraction

OpenStudy (whpalmer4):

Also, you'll have a different factoring for each fraction, as the denominators are different

OpenStudy (anonymous):

no I typed it wrong on the left side it's suppose to be (x-1)(x-6) sorry I can't read my own hand writting tonight

OpenStudy (whpalmer4):

\[(x-1)(x-6) = x^2 - 6x -1x + 6 = x^2 -7x + 6\]Try again? :-)

OpenStudy (whpalmer4):

how about \((x-6)(x+1)\)?

OpenStudy (anonymous):

Yeah I caught that as I was looking it over again :P

OpenStudy (whpalmer4):

\[\frac{(x+8)}{(x+1)(x-6)} + \frac{(x+1)}{(x+1)(x-5)}\]now looks like you need to multiply the left fraction by \(\dfrac{(x-5)}{(x-5)}\); what do you need to do for the right fraction?

OpenStudy (anonymous):

wouldn't the (x+1) on the right cancel top and bottom? right?

OpenStudy (whpalmer4):

It would, if you canceled it, but you don't have to, and don't want to here. You want to get the common denominator for both fractions, then add the numerators

OpenStudy (anonymous):

ok so then I would need to multiply the right side by \[\frac{ (x-6) }{ (x-6) }\]

OpenStudy (whpalmer4):

Yep.

OpenStudy (anonymous):

okay thank you.. I'll tag you if i need any extra help if that is okay?

OpenStudy (anonymous):

this is the final answer I got. \[\frac{ 2x^2-1x-45 }{ (x-5)(x+1)(x-6) }\]

OpenStudy (whpalmer4):

Close but not quite. Check your work for the numerator again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay i rechecked my work... \[\frac{ 2x^2-2x-46 }{ (x-5)(x+1)(x-6) }\]

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Now we're talking! Though you could write the numerator as \(2(x^2-x-23)\). As there's no real definition of what it means to simplify, it's hard to argue with either one.

OpenStudy (anonymous):

okay thank you so much for all the awesome help!!!

OpenStudy (whpalmer4):

You're welcome.